OpenStudy (anonymous):

A disc jockey must play 16 commerical spots during 1 hour of radio show each commerical is either 30 sec or 60 sec long. If the total commerical time is 1 hr and 13 min then how many 30 sec and 60 sec where there? please help I will give a medal!

OpenStudy (anonymous):

I wrote a solution for you

OpenStudy (anonymous):

I can get that sorry......can you please write it here I am sorry

OpenStudy (anonymous):

cant

OpenStudy (anonymous):

Yeah no problem... The key to this solution is this. Set up a system of equations 30x + 60y = 780 x + y = 16 I get this from there are so many 30 second commercials and so many 60 seconds ones that add up to 780 seconds (13min in seconds). 13 minutes is the time spent in commercial during the one hour of radio broadcasting He has to play 16 of these commercials, so we get the second equation. solve the second one in terms of y y = 16-x then plug this into the type 30x + 60(16-x) = 30x - 60x + 960 = 14x + 960 = 780 -180 = -30x x= 6 Plug 6 in for x and you can see y=10. I hope this helps. I offer online tutoring in math and physics Check me out at mathphystut.tripod.com

OpenStudy (anonymous):

so how many in 30 sec how manyi 60 sec

OpenStudy (anonymous):

6 of the 30 second ones and 10 of the 60 second ones

OpenStudy (anonymous):

Thanks, I am really bad at math I am trying to figure this out but I get stuck on some problem!